Respected sir,
[1] A sweet seller declares that he sells sweets at the cost price . However ,he uses a weight of 450 gms instead of 500 gms.His percentage profit is ?
plz help me sir to solve this math problem.
regards.
BINITA

Answer by RAY100(1637)

Breakeven is 500,,,,base
.
sells at 450 units which is (500-450) =50 shy
.
50 /500 = 1/10=.10 = 10 %,,,,profit
.
checking ,,, (.9 ) *500 =450,,, 10 % profit ,,( 1-.9 = .10=10% )